3 Calculating Statistics.doc Ch 13 Calculating Statistics.doc (32 KB) Page 1. Below shown two NBA players scoring points per game Player B 38 29 29 20 36 13 Injured njurec Player A 48 16 4 27 27 35 7 8 a) Which player plays better? Show all the calculations (Tpt)
Answer #1
Game player A Player B |A-mean of A| |B-mean of B|
1 22 28 30.5-22=8.5 20.625-28=7.375
2 48 29 |30.5-48|=17.5 20.625-29=8.375
3 16 29 30.5-16=14.5 8.375
4 22 20 30.5-22=8,5 .625
5 27 36 30.5-27=3.5 15.375
6 27 13 30.5-27=3.5 7.625
7 35 0 |30.5-35|=4.5 20.625
8 27 0 30.5-27=3.5 20.625
TOTAL 224 165 64 88

1) Which player plays better.

To analyse this type of simple data we measure the central tendency of the data and approaches using mean which is the best to calculate the effectiveness of the variables.

Mean of A=( Total points of A) / 8=244 / 8=30.5

Mean of B=(Tolal point of B) /8= 165 / 8=20.625

Hance average point of per game by player A is greater than Player B.

And the maximum point in particular game by player A is greater than Player B.

Hence by the above analysis Player A plays better.

2)Which player performance is more consistent.

Now to check the consistency of the player we calculate the deviation about mean points of the players.

Sum of the deviation of player A=64

Sum of the deviation of player B=88

That is sum of the deviation of the player A is less than player B.

Hence performance of player A more consistent.

Thanks